OpenStudy (anonymous):

Tina bicycles 160 miles at the rate of r mph. The same trip would have taken two hours longer if she had decreased her speed by 4 mph. Find r.

OpenStudy (anonymous):

So we now that 'r' is the speed in mph and 't' is the time in hours Tina took to run the 160 miles. So the ecuation is \[r \times t=160\] On the other hand, going 4 miles slowler would take two hours more 4 mph slowler is \[r -4\] and two hours longer is \[t +2\] So the second ecuation is \[\left( r-4 \right) \times \left( t +2 \right)=160\] So you have two ecuations with two unknown variables To solve it just equalize t like this \[t = 160/r\]and \[t= \left( 160/\left( r-4 \right) \right)-2\] So you have \[160/r=(160/(r-4))-2\] get the r from there

OpenStudy (id21):

Tina bicycles 160/r hours. r-4... Tinas speed if she decreased her first speed by 4 mph (160/r) +2... time, if Tina decreases her speed by 2 mph (r-4)*((160/r)+2)=160... distance is the same, you can make an equation which to solve 160+2r-640/r-8=160 2r-640/r-8=0 r-320/r-4= |*r \[r^2-4r-320=0\] \[r_{1,2}=\frac{4\pm \sqrt{16+1280}}{2}=\frac{4\pm 36}{2}=20;-16\] You are interested in positive values as a positive speed. so \[r=20mph\]

OpenStudy (id21):

* (160/r) +2... time, if Tina decreases her speed by "4" mph